OpenStudy (anonymous):

Solve 20x - 3(7x + 6) ≤ -8x + 16. Write your solution using set builder notaion. Please show all of your work.

OpenStudy (zzr0ck3r):

20x-21x-18<= -8x+16 collect like terms 20x-21x+8x <= 16+18 7x <= 34 x<= 34/7

OpenStudy (anonymous):

Great.....thanks a bunch!!!

OpenStudy (zzr0ck3r):

np, do you understand?

OpenStudy (anonymous):

Sweetie I haven't understood any of this since class started (lol) but I'm trying.